OpenStudy (anonymous):

-3x^2 + 24x need urgent help in finding the maximum

OpenStudy (anonymous):

max value is where \[x=4\] plug it in and you will get your max

OpenStudy (anonymous):

take the derivative, make it equal to zero, find x..that's ur answer

OpenStudy (anonymous):

how did you get tat perfect but i am puzzeled in how you got 4 for x. can you pls write in detail.

OpenStudy (anonymous):

yes. do find the maximum or minimum of a quadratic is the same as finding the second coordinate of the vertex. the first coordinate is always \[-\frac{b}{2a}\] and the second coordinate is whatever you get when you replace x by that value

OpenStudy (anonymous):

in your example \[a=-3,b=24\] so \[\frac{-b}{2a}=4\]

OpenStudy (anonymous):

You need to take the first derivative of the expression. Once you do that, you should get \[-6x+24\] Then you set that equal to 0, and solve for x. so, \[x = 4\] Then plug that x value back into the original equation to get the y value (max value of y). \[-3(4^{2}) + 24(4) = 48\]

OpenStudy (anonymous):

thank you so much!